Rony, who is sixteen years old, is four times as old as his sister. How old will Rony be when he is twice as old as his sister?
Answer by robertb(5830)

Let x = number of years in the future.
If Rony is 16 now, then his sister is 16/4 = years old now.
==> 16 + x = 2(4 + x) ==> 16 + x = 8 + 2x ==> x = 8.
Hence when Rony is 16 + 8 = 24 he will be twice as old as his sister.